A stone is dropped from the top of a tower hits the ground 6 seconds later. How high is the tower?

1
Expert's answer
2021-12-20T10:31:48-0500

g = 9.8 m/s2

V0 = 0

"h=V_0t+\frac{1}{2}gt^2=0+\frac{1}{2}×9.8\frac{m}{s^2}×(6s)^2=176.4\ m"

Answer: 176.4 m